# EEmo-Bench: A Benchmark for Multi-modal Large Language Models on Image Evoked Emotion Assessment

### v3

@@ Abstract.
The furnishing of multi-modal large language models (MLLMs) has led to the emergence of numerous benchmark studies, particularly those evaluating their perception and understanding capabilities. Among these, understanding image-evoked emotions aims to enhance MLLMs’ empathy, with significant applications such as human-machine interaction and advertising recommendations. However, current evaluations of this MLLM capability remain coarse-grained, and a systematic and comprehensive assessment is still lacking. To this end, we introduce EEmo-Bench, a novel benchmark dedicated to the analysis of the evoked emotions in images across diverse content categories. Our core contributions include: 1) Regarding the diversity of the evoked emotions, we adopt an emotion ranking strategy and employ the Valence-Arousal-Dominance (VAD) as emotional attributes for emotional assessment. In line with this methodology, $1,960$ images are collected and manually annotated. 2) We design four tasks to evaluate MLLMs’ ability to capture the evoked emotions by single images and their associated attributes: Perception, Ranking, Description, and Assessment. Additionally, image-pairwise analysis is introduced to investigate the model’s proficiency in performing joint and comparative analysis. In total, we collect $6,773$ question-answer pairs and perform a thorough assessment on $19$ commonly-used MLLMs. The results indicate that while some proprietary and large-scale open-source MLLMs achieve promising overall performance, the analytical capabilities in certain evaluation dimensions remain suboptimal. Our EEmo-Bench paves the path for further research aimed at enhancing the comprehensive perceiving and understanding capabilities of MLLMs concerning image-evoked emotions, which is crucial for machine-centric emotion perception and understanding.
@@
ccs: Human-centered computing Visualization design and evaluation methodsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Understanding image-evoked emotions is crucial for extracting the implicit content from images, especially those without explicit expressive themes. Assessing this capability has wide-ranging applications, particularly in enhancing MLLMs’ empathy and emotional resonance (Spezialetti et al., 2020), which can benefit areas such as human-machine interaction (Minsky, 1988), advertising recommendations (Kaur and Kautish, 2019), public opinion monitoring (Wang and Li, 2015), and image generation (Wu et al., 2023). However, the analysis of image-evoked emotions still faces significant challenges, primarily in two aspects: 1) Viewer responses can vary significantly due to individual divergence, and even a single viewer may experience multiple emotional reactions, emphasizing the need for nuanced and diverse emotional representation. (Peng et al., 2015). 2) Evoked emotions often overlap with the emotions expressed by the image’s subject, a phenomenon that is derived from empathy and emotional resonance (Zhou et al., 2003; Besel and Yuille, 2010), leading to confusion for viewers in determining their primary emotional response.
@@
To bridge these gaps, we introduce EEmo-Bench, a novel benchmark dedicated to systematically evaluating the abilities of MLLMs in perceiving and understanding evoked emotions from images. Through a systematic study of the theoretical frameworks in the IEA field, EEmo-Bench employs seven fundamental emotions based on Ekman’s basic emotions (Ekman, 1982) and incorporates the Valence-Arousal-Dominance (VAD) model (Lang et al., 1997; Warriner et al., 2013) as core emotional attributes, which are widely used in traditional emotion analysis. In terms of the diversity of emotional responses, we apply a ranking strategy for explicit evaluation. The three main evoked emotions for each image are retained and ranked in descending order based on intensity and salience. Unlike emotion distribution methods (Peng et al., 2015), this ranking strategy is better suited for MLLMs, which often struggle to generate balanced distributions (Zhang et al., 2024c). In total, we create a dataset of $1,960$ images with broad content categories, which are manually annotated and strictly scrutinized with the aforementioned emotional attributes.
@@
  • We introduce EEmo-Bench, the first comprehensive benchmark designed to evaluate MLLMs’ ability to perceive and analyze image-evoked emotions. This benchmark includes a meticulously curated dataset of $1,960$ images and $6,773$ manually crafted question-answer (Q&A) pairs.
@@
  • EEmo-Bench incorporates a comprehensive evaluation of $19$ prominent open-source or proprietary MLLMs. The results demonstrate that while Qwen2.5-VL-72B achieves the best overall performance, there remains a substantial gap in comprehensive and fine-grained emotion understanding, providing compelling insights for future advancements.
@@ 2.1. Attributes for Image Emotion analysis
With the growing availability of IEA datasets, the emotion categories involved in them have become more varied. Some datasets use coarse-grained classifications of $6$ to $8$ emotions (Machajdik and Hanbury, 2010; Peng et al., 2015; Lucey et al., 2010; Achlioptas et al., 2021), based on traditional theories such as Ekman’s basic emotions (Ekman, 1982), while others categorize emotions into $26$ or $27$ types for more detailed analysis (Kosti et al., 2017; Demszky et al., 2020; Cowen and Keltner, 2017). In terms of application, the former provides clearer and more representative categorization, making it ideal for the emotion classification of diverse images, while the latter is more meticulous and precise, especially utilized in fields such as aesthetic contexts (Huang et al., 2024b; Zhou et al., 2024a; Mohammad and Kiritchenko, 2018). Despite the clear definition of emotion categories, individual biases in preferences, experiences, and backgrounds can lead to diverse emotional reactions (Joshi et al., 2011), complicating the understanding of image-evoked emotion denotation. In response to this, most studies utilize the dominant emotion in representing the emotion response (You et al., 2016, 2015; Wang et al., 2013; Solli and Lenz, 2010), while some attempt to capture a broader spectrum through emotion distributions (Peng et al., 2015) or descriptive captions (Achlioptas et al., 2023), providing a more comprehensive emotional representation. In addition to categorization, emotional attributes such as valence (You et al., 2016, 2015), arousal (Russell, 1980; Mollahosseini et al., 2019), dominance (Lang et al., 1997; Warriner et al., 2013), and normative significance (Dan-Glauser and Scherer, 2011) have been widely studied, offering a deeper perspective on emotions. Despite the above efforts, there is still a notable gap in existing theories and datasets on applying widely used classical emotional attributes and extending analytical tasks, while both of them are essential for a comprehensive evaluation of emotions.
@@ 2.2. Benchmarks for Image Emotion Analysis
The rapid progress of MLLMs has boosted the development of benchmarks evaluating their perception and understanding capabilities across various visual quality assessment domains, including image quality assessment (IQA) and IAA (Li and Lu, 2024; Fu et al., 2023; Wang et al., 2024b; Wu et al., 2024c; Zhang et al., 2024d; Wu et al., 2024d, b; Huang et al., 2024b; Zhou et al., 2024a). Emotion assessment, as a crucial component for evaluating MLLMs’ high-level understanding abilities, is first introduced in Large Language Model (LLM) benchmarks (Amin et al., 2024; Zhang et al., 2024a), with numerous works emerging, such as EQ-Bench (Paech, 2023) and EmotionQueen (Chen et al., 2024c). However, text-based emotion analysis relies on perceiving emotions through keywords or descriptions, which fundamentally differs from IEA, which involves more complex perceptual mechanisms. Thus, a model’s text-based emotional understanding ability does not accurately reflect its performance in the IEA evaluation. In response to this, early IEA works in IAA apply Q&A tasks designed to probe the dominant emotion and its causes in images, assessing MLLMs’ foundational ability to understand image emotions (Huang et al., 2024b; Zhou et al., 2024a; Huang et al., 2024a). However, there is still a lack of fine-grained and specified image-evoked emotion evaluation benchmarks to thoroughly evaluate MLLMs’ capability to fully perceive and analyze image-evoked emotions categories and attributes. To address this gap, our EEmo-Bench utilizes an emotion ranking strategy to capture diverse emotional experiences and incorporate the VAD model to systematically evaluate MLLMs’ ability to perceive and interpret emotions evoked by images.
@@ 3.3. Source Images Collection
To ensure content diversity, the source images of EEmo-Bench are collected from Flickr, using the seven emotion categories outlined in Sec 3.2 as keywords for retrieving relevant images. Since the original labels often mix expressed emotions and evoked emotions, a manual annotation experiment is conducted to exclusively extract the evoked emotions, with details outlined in Sec. 3.4. In total, we gather $1,960$ images, with $280$ images for each emotion category. To further refine the dataset, content category classification is adopted, including animal, human, stationary object, daily life scene, natural landscape, and abstract/cartoon image.
@@ 3.4. Subjective Annotation
We conduct subjective experiments and recruit $15$ participants for manual annotation in strict compliance with the guidelines outlined in (BT, 2002) and (Sheikh et al., 2006). To accommodate participants unfamiliar with concepts like arousal and dominance, supplementary instructions (like tutorial videos and training sessions) are also provided. (See Supp. A.1 for details.) Each participant is required to complete two tasks for every image: 1) Emotion selection and ranking. Choose up to three emotions from the seven candidate emotions and rank them by intensity from strongest to weakest. 2) VAD rating. Assess the valence, arousal, and dominance levels on the SAM $9$-point scale based on individual emotional response to the image.
After the annotation process, a total of $29,400$ labeled samples are collected. Each sample for the $i$-th image annotated by the $j$-th participant can be denoted as a quadruple ($E,V,A,D$), where $E=[e_{1},e_{2},e_{3}]$ represents a list of emotions, sorted by intensity, with $E\in A_{8}^{3}(e_{ANG},e_{DIS},e_{FEA},e_{JOY},e_{NEU},e_{SAD},e_{SUR},None)$, and $V,A,D$ refers to the valence, arousal, and dominance, respectively. The emotions selected by participants are represented as $e_{EMO}$, where the emotion category can be abbreviated in subscript format. To statistically sort the results of sentiment, we adopt a three-level criterion to calculate the rank of the three main evoked emotions, where the ranking score $S_{EMO}^{i}$ in the $i$-th image is denoted as:
@@ 4.1. Perception
Question types. Two primary question types are engaged. 1) Yes-or-No: To address potential affirmative bias in MLLMs, the ratio of answers is balanced to approximately $1:1$ (Zhang et al., 2024b), ensuring the objectivity of the evaluation. 2) What/How: Designed to capture nuanced emotional perception, emphasizing more precise perception capabilities. While what questions focus on identifying specific emotions and their causes, how questions assess the perception of emotional properties, especially VAD attributes.
@@
Figure 4. A summary of the MLLM’s performance on EEmo-Bench. (a) illustrates the overall performance of the $19$ involved MLLMs. (b) (c) displays comparisons of the top-2 medium-scale open-source MLLMs, large-scale open-source MLLMs, and proprietary MLLMs, throughout different dimensions in perception and description tasks, respectively.
@@
Evaluation process. To address inconsistencies in MLLM outputs, we employ a $5$-round Deepseek-assisted evaluation strategy, which standardizes the outputs and enhances accurate assessment. For efficiency, consistent judgment results across repeated trials are reserved, while manual evaluations are required for the conflicting cases. (Details of the prompt are provided in Supp. A.3.1.)
@@ 4.2. Ranking
(2) $\begin{split}S_{RANk}^{i}&=(\sum_{k=1}^{3}\sum_{j=1}^{3}w_{k}\mathcal{R}(e_{k}^{i},e_{j}^{i}))_{Scaled}\\ &+\mathcal{W}(E_{MLLM}^{i},E_{GT}^{i})\cdot(\mathcal{K}(\mathcal{SP}(E_{MLLM}^{i},E_{GT}^{i}))_{Scaled},\end{split}$
where $\mathcal{R}(e_{k}^{i},e_{j}^{i})$ and $\{w_{k}|k=1,2,3\}$ share the definition in Equ.1, $(\cdot)_{Scaled}$ stands for the scaled function to map the value to the range of $50$, $\mathcal{SP}(\cdot)$ indicates the function to sample the common elements between two emotion lists while preserving their original order, $\mathcal{K}(\cdot)$ represents Kendall’s Tau function, which is used to measure the ranking performance for small sample sizes, and $\mathcal{W}(\cdot)$ denotes a weight function based on the length of the predicted list ($01,\frac{1}{3},$ 0 for the length of $1,2,3$). After that, $S_{RANK}^{i}$ of all images is averaged, representing the final ranking result.
@@ 4.3. Description
We also assess the descriptive abilities of MLLMs. Questions are designed to evaluate comprehensive understanding capabilities based on two categories: 1) Single image analysis. This involves assessing capabilities in detailed analysis, reasoning, direct/indirect emotion evocation, and explaining conflicting emotions. 2) Image pair comparison. This focuses on evaluating the ability to compare emotional properties, including recognizing similarities and differences in emotion ranking and differentiating VAD polarity between two images. To evaluate MLLM performance, we compare their outputs with the golden descriptions using a $5$-round Deepseek-assisted voting as well, measuring Completeness, Accuracy, and Relevance. (Details can be found in Supp. A.3.5.)

### v4

@@ Abstract.
The furnishing of multi-modal large language models (MLLMs) has led to the emergence of numerous benchmark studies, particularly those evaluating their perception and understanding capabilities. Among these, understanding image-evoked emotions aims to enhance MLLMs’ empathy, with significant applications such as human-machine interaction and advertising recommendations. However, current evaluations of this MLLM capability remain coarse-grained, and a systematic and comprehensive assessment is still lacking. To this end, we introduce EEmo-Bench, a novel benchmark dedicated to the analysis of the evoked emotions in images across diverse content categories. Our core contributions include: 1) Regarding the diversity of the evoked emotions, we adopt an emotion ranking strategy and employ the Valence-Arousal-Dominance (VAD) as emotional attributes for emotional assessment. In line with this methodology, 1,9601,9601 , 960 images are collected and manually annotated. 2) We design four tasks to evaluate MLLMs’ ability to capture the evoked emotions by single images and their associated attributes: Perception, Ranking, Description, and Assessment. Additionally, image-pairwise analysis is introduced to investigate the model’s proficiency in performing joint and comparative analysis. In total, we collect 6,7736,7736 , 773 question-answer pairs and perform a thorough assessment on 191919 commonly-used MLLMs. The results indicate that while some proprietary and large-scale open-source MLLMs achieve promising overall performance, the analytical capabilities in certain evaluation dimensions remain suboptimal. Our EEmo-Bench paves the path for further research aimed at enhancing the comprehensive perceiving and understanding capabilities of MLLMs concerning image-evoked emotions, which is crucial for machine-centric emotion perception and understanding. Our code and benchmark datasets are available at https://github.com/workerred/EEmo-Bench.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755777isbn: 979-8-4007-2035-2/2025/10ccs: Human-centered computing Visualization design and evaluation methodsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Understanding image-evoked emotions is crucial for extracting the implicit content from images, especially those without explicit expressive themes. Assessing this capability has wide-ranging applications, particularly in enhancing MLLMs’ empathy and emotional resonance (Spezialetti et al., 2020), which can benefit areas such as human-machine interaction (Minsky, 1988), advertising recommendations (Kaur and Kautish, 2019), public opinion monitoring (Wang and Li, 2015), and image generation (Wu et al., 2023). However, the analysis of image-evoked emotions still faces significant challenges, primarily in two aspects: 1) Viewer responses can vary significantly due to individual divergence, and even a single viewer may experience multiple emotional reactions, emphasizing the need for nuanced and diverse emotional representation. (Peng et al., 2015; Zhao et al., 2022). 2) Evoked emotions often overlap with the emotions expressed by the image’s subject, a phenomenon that is derived from empathy and emotional resonance (Zhou et al., 2003; Besel and Yuille, 2010), leading to confusion for viewers in determining their primary emotional response (Yang et al., 2023).
@@
To bridge these gaps, we introduce EEmo-Bench, a novel benchmark dedicated to systematically evaluating the abilities of MLLMs in perceiving and understanding evoked emotions from images. Through a systematic study of the theoretical frameworks in the IEA field, EEmo-Bench employs seven fundamental emotions based on Ekman’s basic emotions (Ekman, 1982) and incorporates the Valence-Arousal-Dominance (VAD) model (Lang et al., 1997; Warriner et al., 2013) as core emotional attributes, which are widely used in traditional emotion analysis. In terms of the diversity of emotional responses, we apply a ranking strategy for explicit evaluation. The three main evoked emotions for each image are retained and ranked in descending order based on intensity and salience. Unlike emotion distribution methods (Peng et al., 2015), this ranking strategy is better suited for MLLMs, which often struggle to generate balanced distributions (Zhang et al., 2024c). In total, we create a dataset of 1,9601,9601 , 960 images with broad content categories, which are manually annotated and strictly scrutinized with the aforementioned emotional attributes.
@@
  • We introduce EEmo-Bench, the first comprehensive benchmark designed to evaluate MLLMs’ ability to perceive and analyze image-evoked emotions. This benchmark includes a meticulously curated dataset of 1,9601,9601 , 960 images and 6,7736,7736 , 773 manually crafted question-answer (Q&A) pairs.
@@
  • EEmo-Bench incorporates a comprehensive evaluation of 191919 prominent open-source or proprietary MLLMs. The results demonstrate that while Qwen2.5-VL-72B achieves the best overall performance, there remains a substantial gap in comprehensive and fine-grained emotion understanding, providing compelling insights for future advancements.
@@ 2.1. Attributes for Image Emotion analysis
With the growing availability of IEA datasets, the emotion categories involved in them have become more varied. Some datasets use coarse-grained classifications of 666 to 888 emotions (Machajdik and Hanbury, 2010; Peng et al., 2015; Lucey et al., 2010; Achlioptas et al., 2021), based on traditional theories such as Ekman’s basic emotions (Ekman, 1982), while others categorize emotions into 262626 or 272727 types for more detailed analysis (Kosti et al., 2017; Demszky et al., 2020; Cowen and Keltner, 2017). In terms of application, the former provides clearer and more representative categorization, making it ideal for the emotion classification of diverse images, while the latter is more meticulous and precise, especially utilized in fields such as aesthetic contexts (Huang et al., 2024b; Zhou et al., 2024a; Mohammad and Kiritchenko, 2018). Despite the clear definition of emotion categories, individual biases in preferences, experiences, and backgrounds can lead to diverse emotional reactions (Joshi et al., 2011), complicating the understanding of image-evoked emotion denotation. In response to this, most studies utilize the dominant emotion in representing the emotion response (You et al., 2016, 2015; Wang et al., 2013; Solli and Lenz, 2010), while some attempt to capture a broader spectrum through emotion distributions (Peng et al., 2015) or descriptive captions (Achlioptas et al., 2023), providing a more comprehensive emotional representation. In addition to categorization, emotional attributes such as valence (You et al., 2016, 2015), arousal (Russell, 1980; Mollahosseini et al., 2019), dominance (Lang et al., 1997; Warriner et al., 2013), and normative significance (Dan-Glauser and Scherer, 2011) have been widely studied, offering a deeper perspective on emotions. Despite the above efforts, there is still a notable gap in existing theories and datasets on applying widely used classical emotional attributes and extending analytical tasks, while both of them are essential for a comprehensive evaluation of emotions.
@@ 2.2. Benchmarks for Image Emotion Analysis
The rapid progress of MLLMs has boosted the development of benchmarks evaluating their perception and understanding capabilities across various visual quality assessment domains, including image quality assessment (IQA) and IAA (Li and Lu, 2024; Fu et al., 2023; Wang et al., 2024b; Wu et al., 2024c; Zhang et al., 2024d; Wu et al., 2024d, b; Huang et al., 2024b; Zhou et al., 2024a). Emotion assessment, as a crucial component for evaluating MLLMs’ high-level understanding abilities, is first introduced in Large Language Model (LLM) benchmarks (Amin et al., 2024; Zhang et al., 2024a), with numerous works emerging, such as EQ-Bench (Paech, 2023) and EmotionQueen (Chen et al., 2024b). However, text-based emotion analysis relies on perceiving emotions through keywords or descriptions, which fundamentally differs from IEA, which involves more complex perceptual mechanisms. Thus, a model’s text-based emotional understanding ability does not accurately reflect its performance in the IEA evaluation. In response to this, early IEA works in IAA apply Q&A tasks designed to probe the dominant emotion and its causes in images, assessing MLLMs’ foundational ability to understand image emotions (Huang et al., 2024b; Zhou et al., 2024a; Huang et al., 2024a). However, there is still a lack of fine-grained and specified image-evoked emotion evaluation benchmarks to thoroughly evaluate MLLMs’ capability to fully perceive and analyze image-evoked emotions categories and attributes. To address this gap, our EEmo-Bench utilizes an emotion ranking strategy to capture diverse emotional experiences and incorporate the VAD model to systematically evaluate MLLMs’ ability to perceive and interpret emotions evoked by images.
@@ 3.3. Source Images Collection
To ensure content diversity, the source images of EEmo-Bench are collected from Flickr, using the seven emotion categories outlined in Sec 3.2 as keywords for retrieving relevant images. Since the original labels often mix expressed emotions and evoked emotions, a manual annotation experiment is conducted to exclusively extract the evoked emotions, with details outlined in Sec. 3.4. In total, we gather 1,9601,9601 , 960 images, with 280280280 images for each emotion category. To further refine the dataset, content category classification is adopted, including animal, human, stationary object, daily life scene, natural landscape, and abstract/cartoon image.
@@ 3.4. Subjective Annotation
We conduct subjective experiments and recruit 151515 participants for manual annotation in strict compliance with the guidelines outlined in (BT, 2002) and (Sheikh et al., 2006). To accommodate participants unfamiliar with concepts like arousal and dominance, supplementary instructions (like tutorial videos and training sessions) are also provided. (See Supp. A.1 for details.) Each participant is required to complete two tasks for every image: 1) Emotion selection and ranking. Choose up to three emotions from the seven candidate emotions and rank them by intensity from strongest to weakest. 2) VAD rating. Assess the valence, arousal, and dominance levels on the SAM 999-point scale based on individual emotional response to the image.
After the annotation process, a total of 29,40029,40029 , 400 labeled samples are collected. Each sample for the $i$-th image annotated by the $j$-th participant can be denoted as a quadruple ($E,V,A,D$), where $E=[e_{1},e_{2},e_{3}]$ represents a list of emotions, sorted by intensity, with $E\in A_{8}^{3}(e_{ANG},e_{DIS},e_{FEA},e_{JOY},e_{NEU},e_{SAD},e_{SUR},None)$, and $V,A,D$ refers to the valence, arousal, and dominance, respectively. The emotions selected by participants are represented as $e_{EMO}$, where the emotion category can be abbreviated in subscript format. To statistically sort the results of sentiment, we adopt a three-level criterion to calculate the rank of the three main evoked emotions, where the ranking score $S_{EMO}^{i}$ in the $i$-th image is denoted as:
@@ 4.1. Perception
Question types. Two primary question types are engaged. 1) Yes-or-No: To address potential affirmative bias in MLLMs, the ratio of answers is balanced to approximately $:1$ 1 (Zhang et al., 2024b), ensuring the objectivity of the evaluation. 2) What/How: Designed to capture nuanced emotional perception, emphasizing more precise perception capabilities. While what questions focus on identifying specific emotions and their causes, how questions assess the perception of emotional properties, especially VAD attributes.
@@
Figure 4. A summary of the MLLM’s performance on EEmo-Bench. (a) illustrates the overall performance of the 191919 involved MLLMs. (b) (c) displays comparisons of the top-2 medium-scale open-source MLLMs, large-scale open-source MLLMs, and proprietary MLLMs, throughout different dimensions in perception and description tasks, respectively.
@@
Evaluation process. To address inconsistencies in MLLM outputs, we employ a 555-round Deepseek-assisted evaluation strategy, which standardizes the outputs and enhances accurate assessment. For efficiency, consistent judgment results across repeated trials are reserved, while manual evaluations are required for the conflicting cases. (Details of the prompt are provided in Supp. A.3.1.)
@@ 4.2. Ranking
(2) $\begin{split}S_{RANK}^{i}&=(\sum_{k=1}^{3}\sum_{j=1}^{3}w_{k}\mathcal{R}(e_{k}^{i},e_{j}^{i\prime}))_{Scaled}\\ &+\mathcal{W}(E_{MLLM}^{i},E_{GT}^{i})\cdot(\mathcal{K}(\mathcal{SP}(E_{MLLM}^{i},E_{GT}^{i}))_{Scaled},\end{split}$
where $\mathcal{R}(e_{k}^{i},e_{j}^{i})$ and $\{w_{k}|k=1,2,3\}$ share the definition in Equ.1, $(\cdot)_{Scaled}$ stands for the scaled function to map the value to the range of 505050, $\mathcal{SP}(\cdot)$ indicates the function to sample the common elements between two emotion lists while preserving their original order, $\mathcal{K}(\cdot)$ represents Kendall’s Tau function, which is used to measure the ranking performance for small sample sizes, and $\mathcal{W}(\cdot)$ denotes a weight function based on the length of the predicted list ($1,\frac{1}{3},0$ for the length of 1,2,31,2,31 , 2 , 3). After that, $S_{RANK}^{i}$ of all images is averaged, representing the final ranking result.
@@ 4.3. Description
We also assess the descriptive abilities of MLLMs. Questions are designed to evaluate comprehensive understanding capabilities based on two categories: 1) Single image analysis. This involves assessing capabilities in detailed analysis, reasoning, direct/indirect emotion evocation, and explaining conflicting emotions. 2) Image pair comparison. This focuses on evaluating the ability to compare emotional properties, including recognizing similarities and differences in emotion ranking and differentiating VAD polarity between two images. To evaluate MLLM performance, we compare their outputs with the golden descriptions using a 555-round Deepseek-assisted voting as well, measuring Completeness, Accuracy, and Relevance. (Details can be found in Supp. A.3.5.)